# Representative electricity price profiles for European day-ahead and intraday spot markets

### v1

@@ 2.2 ID price profile
We proceed with generating an average ID profile by first exploring the relationship between DA and ID prices, followed by the construction of an ID profile based on a DA reference. Rahimiyan & Baringo, (2016) assessed the correlation between DA and ID prices, pinpointing uncorrelated DA price and market deviation (also referred to as DA-ID price deviation or residual price). As noticed by Germscheid et al., (2022), the ID price oscillates around the DA price in a well-defined, season-independent harmonic pattern with dominant frequencies at $0.5\,\mathrm{h^{-1}}$ and close to $1\,\mathrm{h^{-1}}$. Consequently, representative ID profiles may be constructed by first subtracting ID and DA data and then using this market deviation to construct a harmonic deviation profile. Here, we build on these findings and introduce some further refinements. We consider the $\text{ID}_{\text{3}}$ price index, which is the volume-weighted average price of all trades that took place from three hours to 30 min before delivery (EPEX,, 2023).
We first calculate the average market deviation of ID around the DA profile for each 15 min interval of the historical data ($\overline{\Delta}_{q}^{ID}$) similar to Germscheid et al., (2022):
$\begin{split}\Delta_{i,q}^{ID}&=ID_{i,q}-DA_{i,\lceil q/4\rceil}\,,\\ \overline{\Delta}_{q}^{ID}&=\frac{1}{N}\sum_{i=1}^{N}\Delta_{i,q}^{ID}\,,\end{split}$ (3)
@@
Therein, we observe the deviation of an hourly pattern as indicated by a Fourier analysis (Germscheid et al.,, 2022). Notably, large jumps in the price deviation occur at full hours, typically involving a change of sign. Moreover, we observe phase shifts at hours 3, 7, 13, and 19, where the inner hour trend switches between continuously increasing and decreasing. Next, a straightforward approach for constructing an ID profile would be (Germscheid et al.,, 2022):
@@
such that the cumulative prices of the benchmark DA and ID profiles will be identical. Figure 4(a) reveals that the daily cumulative deviation scaled by the daily cumulative DA price in 2023 is approximately symmetrically distributed around zero, which supports Eq. (5). Moreover, even the hourly integral of the ID to DA price deviations scaled by the hourly cumulative DA price appears to be symmetric around zero mean, Fig. 4(b). This observation is consistent with the dominant harmonic frequencies at $0.5\,\mathrm{h^{-1}}$ and $1\,\mathrm{h^{-1}}$ mentioned above. We found similar distributions for the years 2019 to 2022 (figures omitted for brevity). The distributions, although approximately zero-mean, exhibit significant deviations. Hence, they only valid to first order.
@@
Because the distribution in Fig. 4(a) is, however, not perfectly Gaussian and exhibits non-zero mean (2.23 €/MWh, also see Fig. 3), Eq. (3) does not satisfy the comparability requirement (Eq. (5)). Hence, we modify the approach and extend Eq. (4) by an offset correction term to obtain:
@@
We notice that the correction approach may be further refined based on the zero-mean distribution in Fig. 4(b). However, we decide against such an additional refinement to keep the method simple. Next, similar to $\beta$ for the DA profile, we scale $\tilde{\Delta}^{ID}_{q}$ by a parameter $\gamma>0$ to match the year-average of the ID standard deviations:
@@
wherein the benchmark ID profile is calculated as:
@@
The nominal parameter value $\gamma$ is specified by solving the implicit nonlinear equation with $\beta$ fixed:
@@
$\begin{split}Q_{3.4}\big{(}\{\sigma(ID_{i,1},...,ID_{i,96})\}_{i=1}^{N}\big{)}&=\sigma(\overline{ID}_{1}^{\gamma},...,\overline{ID}_{96}^{\gamma})\\ &=\sqrt{\frac{1}{96}\sum_{q=1}^{96}(\beta(\overline{DA}_{\lceil q/4\rceil}-\overline{DA})+\gamma\tilde{\Delta}_{q}^{ID})^{2}}\,,\end{split}$ (9)
@@
Figure 5 illustrates the final single-day nominal benchmark DA and ID profile based on full year 2023 data and for $\beta=1.47$ and $\gamma=1.91$. For the DA profile we note that prices exhibit a daily minimum, coinciding with peak photovoltaic generation around noon (Schäfer et al.,, 2020; Kiesel & Paraschiv,, 2017). Prices start rising around 6:00, when the workday begins and start falling around 18:00 when the workday ends. Two price peaks around 7:00 and 18:00 denote a high electricity demand meeting a low renewable energy supply from photovoltaics (Knittel & Roberts,, 2005). On the other hand, ID prices drop subhourly between 07:00 to 13:00 and 19:00 to 03:00 and rise in the meantime. Namely, while the sun is rising (from 07:00 to 13:00), offer is higher than the hourly average demand and thus price is lower in the last quarter of the hour. During afternoon hours (from 13:00 to 19:00), the offer is higher in the first quarter of the hour. Further, during offpeak hours (from 19:00 to 07:00) the oscillation of the average ID price follows the demand of power-intensive industry (decreasing steps), and the production design of the fossil power plants (increasing steps) for minimum production regulations (Kiesel & Paraschiv,, 2017).
Fig. 5: Nominal ($\beta=1.47$, $\gamma=1.91$) single-day DA and ID price profile for year 2023 (EPEX,, 2023).
@@ 3.1 DA price profile
Commonly, the time horizon of planning and scheduling spans more than one day of operation. In some cases, a periodic scenario is sufficient, wherein the price profile is duplicated for multiple consecutive days. However, if distinct prices are desired, the procedure from the previous section can be extended to price scenarios comprising multiple days. Here, we construct such a price profile for a one-week time window. The choice of this horizon length aligns with the significant DA price frequencies pinpointed by Germscheid et al., (2022), allowing us to capture and average over weekly patterns. However, we do not recommend to apply the averaging method to construct time frames longer than a week. In that case, the averaging would involve either statistically too small historical data sets (for $N=365$) or data sets spanning more than a year and thus including long-term changes of market dynamics and policy. However, one remedy to construct a profile spanning multiple weeks could be the periodic repetition of the averaged profile.
Following the approach presented in Section 2.1, we first average the historical DA prices for every hour $16824d+k=1,2,...,$ 168 of the week after splitting the available data in an integer number of full weeks $W$ (Monday to Sunday), so that $7\cdot W\leq N$:
@@
where $d=0,1,...,6$ indicates the day of the week $w=1,...,W$ (e.g, Monday corresponds to 0), $k=1,2,...,24$ denotes the hour of the day and $24d+k$ is the hour of the constructed single-week profile. Similar to Section 2.1 we use the hourly data of 2023, thus considering $W=52$.
@@
For the construction of a nominal benchmark profile, $\beta$ is selected such that the standard deviation of the constructed single-week profile is equal to the average standard deviation of all weekly profiles of the historical data:
@@ 3.2 ID price profile
We extend our methodology on extracting a representative ID profile directly following the analysis in Section 2.2 to compose the ID profile over a one-week horizon by superposing the average ID-DA price deviation with the average DA price over a single-week period. Therefore, we first calculate the average market deviation of the historical data:
$\begin{split}\overline{\Delta}_{96d+q}^{ID}&=\frac{1}{W}\sum_{w=1}^{W}\big{(}ID_{(d+1)+7(w-1),q}-DA_{(d+1)+7(w-1),\lceil q/4\rceil}\big{)}\,,\\ \end{split}$
where $q=1,...,96$ is the quarter hour index, $ID_{(d+1)+7(w-1),q}$ is the ID price on week $w$, day of the week $d$ and quarter hour $q$.
@@
Considering the historical data of EPEX Spot for 2023, and 52 full weeks (from Monday to Sunday), Eq. (12) gives $\gamma=1.60$, and Eq. (13) gives $\gamma=2.32$. The final single-week nominal benchmark DA and ID profiles for the full year 2023 data ($\beta=1.58$,$\gamma=1.60$) are shown in Fig. 6.
Figure 6 is consistent with the observations of Germscheid et al., (2023). Lowest prices are reached during the weekend associated with a lower demand. Accordingly, the price peak in the weekend mornings is less significant than the peak in the evening. As reported by Guthrie & Videbeck, (2002), during weekdays the daytime off-peak and evening peak periods exhibit distinct market behaviors. However, during the weekend these periods operate as a unified market. Additionally, the absence of morning peaks in the weekends results from morning prices correlated with the rest prices of the weekend. The highest DA prices are reached on Monday and Thursday evening, and the highest ID prices on Monday evening.
@@
4   Exemplary scenarios and comparison to historical data
We apply the proposed method to calculate DA and ID price profiles based on the full year 2023 historical data (01.01.2023 to 31.12.2023). We construct profiles following Sections 2 and 3 with three variance scaling cases each:
@@ 4   Exemplary scenarios and comparison to historical data
    i) Nominal ($\beta=1.47$, $\gamma=1.91$), ii) Moderate ($\beta=1.0$, $\gamma=1.0$), iii) Extreme ($\beta=1.85$, $\gamma=1.77$),
@@
    i) Nominal ($\beta=1.58$, $\gamma=1.6$), ii) Moderate ($\beta=1.0$, $\gamma=1.0$), iii) Extreme ($\beta=1.76$, $\gamma=2.32$).
We show the resulting single-day nominal profiles in Fig. 5 and provide the moderate and extreme profile in the SI. Similarly, the single-week nominal profiles are shown in Fig. 6 and the moderate and extreme profiles can be found in the SI. We furthermore provide the profiles in CSV format via Git (Link) as well as in table format in the SI. We highlight that the Moderate data may be used as a basis to construct price profiles for arbitrary scaling options ($\beta$, $\gamma$).
@@
Fig. 8: Moderate single-day and corresponding best fit of historical profile for year 2023.
(a) DA market
(b) ID market
Fig. 9: Nominal single-week price profiles compared to the best fit historical week in 2023.
For the sake of comparison, we complement the constructed profiles by historical profiles that closely match each profile. The historical profiles are retrieved by minimizing the mean absolute deviation between the historical and a constructed scenario for DA and ID simultaneously. In the single-day case, we obtain:
$\min_{i}\Big{(}\sum_{k=1}^{24}|\overline{DA}_{k}^{\beta}-DA_{i,k}|+\frac{1}{4}\sum_{q=1}^{96}|\overline{ID}_{q}^{\gamma}-ID_{i,q}|\,.\Big{)}$ (14)
A single-week historical profile is identified analogously. Figure 7 contrasts the nominal single-day profiles and the historical day with the best fit. While the inner hour variance of the nominal ID profile is more pronounced during early and late hours, all profiles follow similar trends and are overall comparable. Moreover, inserting Eq. (2) into Eq. (14) reveals that the historic profile in 7 indeed reflects the day which deviates the least from the rest historical days, i.e., the cluster center in this sense. Next, Fig. 8 depicts the moderate profiles and the historical best fit. Since here all single-week profiles have the same historical week as the best fit, we only show the plot for the nominal profile in Fig. 9. This approach is equivalent to determining the cluster center of a single cluster formed by all available time series. We conclude that the moderate and nominal profile represent the historical data closely. For the nominal profile, the fitted historical day merely serves for validation, whereas for the moderate profile the historical day corresponds to the most representative day in the historical data set.
Next, we discuss a limitation of our approach. Compared to the close alignment of the moderate and nominal profiles with historical data, solving the fitting problem in Eq. (14) for the extreme profile did not lead to satisfactory results. This is explained by the extreme deviations or outliers observed in the historical data, especially in the ID price. These outliers, although leading to a high variance, which is considered for when composing the extreme profiles, exhibit different patterns than the averaged ones of the constructed (scaled) profiles. Nevertheless, when fitting historical DA and ID profiles separately rather than jointly, we were able to find similar historical profiles (see figures in SI). We conclude that while a more extreme variance scaling is generally valid, an alternative scenario generation method should be consulted if extreme profiles with abnormal patterns are sought.
In addition to the qualitative assessment of the profiles by the figures, we provide a quantitative statistical evaluation in Tables 4 to 4. In particular, the tables confirm that the standard deviation of the constructed nominal DA and ID profiles and the corresponding historical data match as intended by nominal scaling. Moreover, while the mean of the DA profiles matches the historical data, there is a difference for the ID profiles due to the correction of the integral zero-mean deviation. Consequently, the 24 h price integrals of the corresponding DA and ID profiles are identical (compare last row in Tables 4 and 4 as well as Tables 4 and 4).
@@
🟊 The number of histogram bins is calculated with the Scott rule (Scott,, 1979)
@@
Nominal Moderate Extreme historical
@@
Nominal Moderate Extreme historical
@@
Nominal Moderate Extreme historical
@@
Nominal Moderate Extreme historical

### v2

@@ 3.2 ID price profile
We proceed with generating an average ID profile by first exploring the relationship between DA and ID prices, followed by the construction of an ID profile based on a DA reference. Rahimiyan and Baringo (2016) assessed the correlation between DA and ID prices, pinpointing uncorrelated DA price and market deviation (also referred to as DA-ID price deviation or residual price). As noticed by Germscheid et al. (2022), the ID price oscillates around the DA price in a well-defined, season-independent harmonic pattern with dominant frequencies at $0.5\,\mathrm{h^{-1}}$ and close to $1\,\mathrm{h^{-1}}$. We perform a similar analysis for the DA prices and the ID-DA deviation of 2023 (frequency spectra provided in the SI) that confirms the observations of Germscheid et al. (2022). Consequently, ID profiles may be constructed by first subtracting ID and DA data and then using this market deviation to construct a harmonic deviation profile. Here, we build on these findings and introduce some further refinements. We consider the $\text{ID}_{\text{3}}$ price index, which is the volume-weighted average price of all trades that took place within the last three hours before delivery (EPEX, 2023).
We first calculate the average market deviation of ID around the DA profile for each 15 min interval of the historical data ($\overline{\Delta}_{q}^{ID}$) similar to Germscheid et al. (2022):
@@
where q=1,,96q=1,...,96 is the quarter hour index, $ID_{i,q}$ is the ID price on day $i$ and quarter hour $q$, and $\lceil\,\cdot\,\rceil$ denotes the ceiling operation utilized to translate from quarter hour index $q$ to corresponding hour index $k$. Fig. 3 shows an exemplary profile of $\overline{\Delta}_{q}^{ID}$ based on the full year 2023.
@@
Therein, we observe the deviation of an hourly pattern as indicated by a Fourier analysis (Germscheid et al., 2022). Notably, large jumps in the price deviation occur at full hours, typically involving a change of sign. Moreover, we observe phase shifts at hours 3, 7, 13, and 19, where the inner hour trend switches between continuously increasing and decreasing. Next, a straightforward approach for constructing an ID profile would be:
@@
such that the cumulative prices of the DA and ID profiles will be identical. Figure 4(a) reveals that the daily cumulative deviation scaled by the daily cumulative DA price in 2023 is approximately symmetrically distributed around zero, which supports Eq. (5). Moreover, even the hourly integral of the ID to DA price deviations scaled by the hourly cumulative DA price appears to be symmetric around zero mean, Fig. 4(b). This observation is consistent with the dominant harmonic frequencies at $0.5\,\mathrm{h^{-1}}$ and $1\,\mathrm{h^{-1}}$ mentioned above. We found similar distributions for the years 2019 to 2022 (figures omitted for brevity). The distributions, although approximately zero-mean, exhibit significant deviations. Hence, they only valid to first order.
@@
Because the distribution in Fig. 4(a) is, however, not perfectly Gaussian and exhibits non-zero mean (2.23 €/MWh, also see Fig. 3), Eq. (3) does not satisfy the consistency requirement (Eq. (5)). Hence, we modify the approach and extend Eq. (4) by an offset correction term to obtain:
@@
We notice that the correction approach may be further refined based on the zero-mean distribution in Fig. 4(b). However, we decide against such an additional refinement to keep the method simple. We remark that analysis of the years 2019 to 2022 revealed both positive and negative correction terms of smaller orders of magnitude compared to 2023 (2.23 €/MWh). This indicates that the correction assumption is generally valid and introduces minimal changes to the averaged historical data. Next, similar to $\beta$ for the DA profile, we scale $\tilde{\Delta}^{ID}_{q}$ by a parameter $\gamma>0$ to match the year-average of the ID standard deviations:
@@
wherein the constructed ID profile is calculated as:
@@
The parameter value $\gamma$ corresponding to the Nominal profile is specified by solving the implicit nonlinear equation with $\beta$ fixed:
@@
[…]
@@
Figure 5 illustrates the final single-day Nominal ready-to-use DA and ID profile based on full year 2023 data and for $\beta=1.47$ and $\gamma=1.91$. For the DA profile we note that prices exhibit a daily minimum, coinciding with peak photovoltaic generation around noon (Schäfer et al., 2020; Kiesel and Paraschiv, 2017). Prices start rising around 6:00, when the workday begins and start falling around 18:00 when the workday ends. Two price peaks around 7:00 and 18:00 denote a high electricity demand meeting a low renewable energy supply from photovoltaics (Knittel and Roberts, 2005). On the other hand, ID prices drop subhourly between 07:00 to 13:00 and 19:00 to 03:00 and rise in the meantime. Namely, while the sun is rising (from 07:00 to 13:00), offer is higher than the hourly average demand and thus price is lower in the last quarter of the hour. During afternoon hours (from 13:00 to 19:00), the offer is higher in the first quarter of the hour. Further, during offpeak hours (from 19:00 to 07:00) the oscillation of the average ID price follows the demand of power-intensive industry (decreasing steps), and the production design of the fossil power plants (increasing steps) for minimum production regulations (Kiesel and Paraschiv, 2017).
Fig. 5: Nominal ($\beta=1.47$, $\gamma=1.91$) single-day DA and ID price profile for year 2023 (EPEX, 2023).
@@ 4  Single-week price profile
In this section, we adust the methodology presented in the previous sections to generate single-week DA (Section 4.1) and ID (Section 4.2) price profiles from the full-year time-series data set.
@@ 4.1 DA price profile
Commonly, the time horizon of planning and scheduling spans more than one day of operation. In some cases, a periodic scenario is sufficient, wherein the price profile is duplicated for multiple consecutive days. However, if distinct prices are desired, the procedure from the previous section can be extended to price scenarios comprising multiple days. Here, we construct such a price profile for a one-week time window. The choice of this horizon length aligns with the significant DA price frequencies pinpointed by Germscheid et al. (2022), allowing us to capture and average over weekly patterns. However, we do not recommend to apply the averaging method to construct time frames longer than a week. In that case, the averaging would involve either statistically too small historical data sets (for $N=365$) or data sets spanning more than a year and thus including long-term changes of market dynamics and policy. However, one remedy to construct a profile spanning multiple weeks could be the periodic repetition of the averaged profile.
Following the approach presented in Section 3.1, we first average the historical DA prices for every hour 24d+k=1,2,,16824d+k=1,2,...,168 of the week after splitting the available data in an integer number of full weeks $W$ (Monday to Sunday), so that $7\cdot W\leq N$:
@@
where d=0,1,,6d=0,1,...,6 indicates the day of the week w=1,,Ww=1,...,W (e.g, Monday corresponds to 0), k=1,2,,24k=1,2,...,24 denotes the hour of the day and $24d+k$ is the hour of the constructed single-week profile. Similar to Section 3.1 we use the hourly data of 2023, thus considering $W=52$.
@@
For the construction of a Nominal profile, $\beta$ is selected such that the standard deviation of the constructed single-week profile is equal to the average standard deviation of all weekly profiles of the historical data:
@@ 4.2 ID price profile
We extend our methodology on extracting a practical ID profile directly following the analysis in Section 3.2 to compose the ID profile over a one-week horizon by superposing the average ID-DA price deviation with the average DA price over a single-week period. Therefore, we first calculate the average market deviation of the historical data:
[…]
where q=1,,96q=1,...,96 is the quarter hour index, $ID_{(d+1)+7(w-1),q}$ is the ID price on week $w$, day of the week $d$ and quarter hour $q$.
@@
Considering the historical data of EPEX Spot for 2023, and 52 full weeks (from Monday to Sunday), Eq. (12) gives $\gamma=1.60$, and Eq. (13) gives $\gamma=2.32$. The final single-week Nominal ready-to-use DA and ID profiles for the full year 2023 data ($\beta=1.58$$\gamma=1.60$) are shown in Fig. 6.
Figure 6 is consistent with the observations of Germscheid et al. (2023). Lowest prices are reached during the weekend associated with a lower demand. Accordingly, the price peak in the weekend mornings is less significant than the peak in the evening. As reported by Guthrie and Videbeck (2002), during weekdays the daytime off-peak and evening peak periods exhibit distinct market behaviors. However, during the weekend these periods operate as a unified market. Additionally, the absence of morning peaks in the weekends results from morning prices correlated with the rest prices of the weekend. The highest DA prices are reached on Monday and Thursday evening, and the highest ID prices on Monday evening.
@@
To capture seasonal price patterns, the approach could be extended by averaging over a 3-month period. However, we consider the data from a single calendar year too limited for this technique. At the same time, using data from 2022 and 2023 is not feasible due to significant deviations caused by the 2022 energy crisis.
5  Exemplary price profiles
In this section, we present the price scenarios derived from our methodology application (Section 5.1) and compared the resulting profiles to historical time-series (Section 5.2).
5.1 Scenarios
We apply the proposed method to calculate DA and ID price profiles based on the full year 2023 historical data (01.01.2023 to 31.12.2023). We construct profiles following Sections 3 and 4 with three variance scaling cases each:
@@ 5.1 Scenarios
    i) Nominal ($\beta=1.47$, $\gamma=1.91$), ii) Unscaled ($\beta=1.0$, $\gamma=1.0$), iii) Extreme ($\beta=1.85$, $\gamma=1.77$),
@@
    i) Nominal ($\beta=1.58$, $\gamma=1.6$), ii) Unscaled ($\beta=1.0$, $\gamma=1.0$), iii) Extreme ($\beta=1.76$, $\gamma=2.32$).
We show the resulting single-day Nominal profiles in Fig. 5 and provide the Unscaled and Extreme profile in the SI. Similarly, the single-week Nominal profiles are shown in Fig. 6 and the Unscaled and Extreme profiles can be found in the SI. We furthermore provide the profiles in CSV format via Git (Link) as well as in table format in the SI. We highlight that the Unscaled data may be used as a basis to construct price profiles for arbitrary scaling options ($\beta$, $\gamma$).
@@
Fig. 8: Unscaled single-day and corresponding best fit of historical profile for year 2023.
Fig. 9: Nominal DA single-week price profiles compared to the best fit historical week in 2023.
We additionally provide a quantitative statistical evaluation of the constructed profiles in Tables 5 to 5. In particular, the tables confirm that the standard deviation of the Nominal DA and ID profiles and the corresponding historical data match as intended by nominal scaling. Moreover, while the mean of the DA profiles matches the historical data, there is a difference for the ID profiles due to the correction of the integral zero-mean deviation. Consequently, the 24 h price integrals of the corresponding DA and ID profiles are identical (compare last row in Tables 5 and 5 as well as Tables 5 and 5).
@@
🟊 The number of histogram bins is calculated with the Scott rule (Scott, 1979)
@@
Nominal Unscaled Extreme historical
@@
Nominal Unscaled Extreme historical
@@
Nominal Unscaled Extreme historical
@@
Nominal Unscaled Extreme historical